OpenStudy (anonymous):

A boy can accelerate at 1.00 m/s 2 over a short distance. If the were to take an energy pill and suddenly have the ability to accelerate at 5.6 m/s 2, then how would his new energy pill force compare to his earlier force? If the boy's earlier force was 45 N, what is the size of his energy pill force?

OpenStudy (anonymous):

I need to get me some of these energy pills. :-) First, let's recall Newton's Second Law. Mathematically, it is modeled as\[F = m \cdot a\]where \(m\) is the boy's mass and \(a\) is the boy's acceleration. What can we observe about the boy's force if we hold mass constant but increase acceleration?

OpenStudy (kunal):

now the force is about 252N but how to calculate energy??

OpenStudy (stormfire1):

As eashmore stated, start with \[F=ma\]You're given his initial force and acceleration: \[45N=m * 1.0 m/s ^2 \]Therefore his mass must be 45 kg. To find the force after taking the pill plug in what you now know: \[F = 45kg * 5.6 m/s^2 = 252N\]Now just look at the ratio: \[\frac{252N}{45N} = 5.6\] Therefore his force with the pill was 5.6 times larger than the force without it. You can also figure this out more easily by noting that if mass is constant force is directly proportional to the acceleration...and just skip to the final ratio above.